You're evaluating the performance of your pension fund. You invested $100 initially, which grew to $104 after 4 months, and then to $108 after another 5 months.

What was your HPR during the first 4 months?

What was your HPR during the next 5 months?
What was your total HPR over the 9 months?
1. Youre evaluating the performance of your pension fund. You invested $100 initially, vhich grew to $104 after 4 months, and then to $108 after another 5 months. a. What was your HPR during the first 4 months? b. What was your HPR during the next 5 months? c. What was your total HPR over the 9 months?
Answer #1

Holding period return is return on investment based on the holding period of that particular investment, below formula is used to calculate the HPR.

Formula to calculate HPR for single period = Earnings+Asset appreciation
Initial investment

A) HPR during the first 4 months is = (104-100)/100*100= 4%

HPR per month is 4%/ 4 months= 1% per month.

B) HPR for the next 5 months is = (108-104)/104*100= 3.85%

HPR per month = 3.85%/5 months =0.769%

Formula to calculate HPR annually, Quarterly or monthly is = {(1+R1)X(1+R2)X…..X(1+Rn)}-1

Where R=% of return n per period

n= Number of periods

C)HPR over the 9 months =

=((1+0.01)*(1+0.01)*(1+0.01)*(1+0.01)*(1+0.000769)*(1+0.000769)*(1+0.000769)*(1+0.000769)*(1+0.000769))-1= 4%